OpenStudy (roadjester):

Find dy/dx and d^2y/dx^2. For which values of t is the curve concave upward? x=t-e^t y=t+e^(-t)

OpenStudy (roadjester):

\[x=t-e^t\] \[y=t+e^{-t}\]

OpenStudy (anonymous):

I have this answer

OpenStudy (roadjester):

Help please @Ket-kat

OpenStudy (anonymous):

x + y = e −t + e t = 2cosh(t). Therefore t = arccosh((x + y)/2)

OpenStudy (roadjester):

uh @Ket-kat , what did you do?

OpenStudy (anonymous):

http://www.aerostudents.com/files/calculus/solutions/calculus_solutions_10_2.pdf I think this will make more sense for you; they are similar questions so you can see how to work them out. If you don't get it, message me :)